A laser shines through a pinhole and it makes a diffraction pattern on a screen behind. The area of the circular hole is reduced by a factor of 2. What can you say about the new intensity? A. less than 1/2 of original intensity B. 1/2 the original intensity C. same as original intensity D. greater than the original intensity
Answer #1

Solution) The intensity I = P/A

So if Area is reduced by factor of 2 then it's intensity changes by

I' = P/(A/2) = 2P/A

I' = 2I

Therefore the new intensity is greater than original intensity .

Option D